OpenStudy (anonymous):

A plane landing at PDX has to descend 1650ft over a diagonal distance of 9700ft. What is the angle of depression of the plane? help would be nicce<33

OpenStudy (anonymous):

\[\frac{\text{ArcSin}[1650/9711]}{\pi }180 = 9.78261 \text{ Degrees} \]

OpenStudy (anonymous):

oh gotcha thanks!

OpenStudy (anonymous):

Sorry. Entered 9711 and it should have been 9700. Not much difference in the angle. \[\frac{\text{ArcSin}[1650/9700]}{\pi }180 = 9.79381 \text{ degrees} \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

no prob<3

OpenStudy (anonymous):

myininaya, Here's what I got: \[\text{Cos}\left[\text{ArcSec}\left[\frac{5}{4}\right]+\text{ArcSin}\left[\frac{1}{3}\right]\right]=-\frac{1}{5}+\frac{8 \sqrt{2}}{15}\]